OpenStudy (anonymous):

Solve the equation 5^x = 6.2. Round your answer to the nearest hundredth. A. x ≈ 1.13 B. x ≈ 2.59 C. x ≈ 1.45 D. x ≈ 3.49

OpenStudy (anonymous):

do you have a guess?

OpenStudy (anonymous):

i have a guess by the change of base formula \[5^x=6.2\iff x=\frac{\ln(6.2)}{\ln(5)}\] and a calculator

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

yw